OpenStudy (anonymous):

can some one help me solve this problem 6=1/5x+11

OpenStudy (lucidiousxiv):

I take it you're solving for X?

OpenStudy (anonymous):

yes

OpenStudy (lucidiousxiv):

X would be 25.

OpenStudy (lucidiousxiv):

1 Multiply both sides by 5 30=x+55 2 Subtract 55 from both sides 30−55=x 3 Simplify 30−55 to −25 −25=x 4 Switch sides x=−25

OpenStudy (rabya!):